jonesc21p52.problem Due this Thursday, Apr 12 at 11:59 pm (EDT) 4.30×101 ?, an inductor L-6.60x10-2 H, and a battery of emf-6 Volts has a time constant Q-The circuit is closed at time t-o. what is the time constant of this (c21 p52) A series circuit of a resistor R circuit? 1.53x10-3 s ou are correct. r receipt no. is 161-5520 Previous Tries What is the value of the current in the circuit when t 7.67E-4 s? Submit Answer I Tries 3/10 Previous Tries When t-3.07E-3 s? Submit Answer Tries 0/10 When t 6.14E-3 S? Submit Answer Tries o/10
Answer #1

(b) Current I in the series R-L circuit is -

i(t) = V*[1 - e^-(R/L)t] / R

t = 7.67 x 10^-4 s

i(t) = 6*[1 - e^-(43/0.066)*(7.67x10^-4)] / 43 = 0.05487 A = 54.87 mA

(c) t = 3.07 x 10^-3 s

i(t) = 6*[1 - e^-(43/0.066)*(3.07x10^-3)] / 43 = 0.12065 A = 120.65 mA

(d) t = 6.14 x 10^-3 s

i(t) = 6*[1 - e^-(43/0.066)*(6.14x10^-3)] / 43 = 0.1370 A = 137.0 mA